OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (anonymous):

The first thing we want to do is to find the slope between these two points. The equation for the slope, m, is: example m=rise/run=(y2-y1)/(x2-x1). Let the first point (-1,0)=(x1,y1) and the second point (1,2)=(x2,y2), then m=(2-0)/(1-(-1))=2/2=1, so the slope between these two points is 1. Now, we can write the equation of the line in point-slope form, where m is the slope: y-y1=m(x-x1) y-0=1(x-(-1)) y-0=1(x+1) or y=1(x+1), which is the equation of the line in point-slope form. Note: Instead of (x1,y1), you could also use (x2,y2). Your point-slope form equation will then look like: y-2=1(x-1), which will still give you the same equation for the slope-intercept form once you rearrange the equation. The equation for slope-intercept form is, where m is the slope and b is the y-intercept: y=mx+b

OpenStudy (anonymous):

so to find the slope it would be m=(5-4)(2-(-1) Since that is double negative it becomes positive and the equation would then be m=(5-4)(2+1)

OpenStudy (anonymous):

rise over run you need to divide to find the slope

OpenStudy (anonymous):

1/3x then find your y intercept

OpenStudy (anonymous):

pretty much d:

OpenStudy (anonymous):

m=(5-4)/(2+1) so the slope is 3

OpenStudy (anonymous):

yes opps i made a mistake

OpenStudy (anonymous):

no it came out to 1/3 but i just changed it to 3

OpenStudy (lopezking1):

(2 , 5) (–1, –4) ------- 3 , 9 ; slope = 9/3 = 3 the equation of the line is: y = mx -mPx + Py , where (Px,Py) is any point that is given

OpenStudy (anonymous):

what is the P for?

OpenStudy (lopezking1):

any point that is given lol

OpenStudy (anonymous):

lol oh okay

OpenStudy (anonymous):

okay so i have the slope now what?

OpenStudy (lopezking1):

and m=slope y = 3x -3(2)+5 y = 3x -6+5 and y equals what

OpenStudy (anonymous):

2

OpenStudy (lopezking1):

close but no try again -6+5= what y=mx+b y=3x- (notice the negative sign thus your answer will be negative)

OpenStudy (anonymous):

y=3x-1!!!

OpenStudy (lopezking1):

Yess lol :D since we used both points to form the line; it really doesnt matter which one we use to calibrate the equation with

OpenStudy (anonymous):

Omg thank you guys so much idk why it was so hard for me to figure out!

OpenStudy (lopezking1):

Lol because its math and math sucks :)

OpenStudy (anonymous):

lol its not math so much it just Algebra!!